I submitted this question yesterday, but left out some vital information. Sorry! :-( Could I just blame it on age? Anyway, here we go again:
Three darts are thrown at the target shown below. Assume that each of the darts lands within one of the rings, or within the bull's eye. How many different point totals are possible?
Three rings- Outer ring - 3 points
Center ring - 4 points
Bull's eye - 9 points
Thanks for your patience.

Answer by edjones(8007)

3*3=9
3+3+4=10
3+4+4=11
3*4=12
9+3+3=15
9+3+4=16
9+4+4=17
9+9+3=21
9+9+4=22
9*3=27